If magnitude of vector a is 2 and magnitude of vector b is 5 and magnitude of a × b is 8 then what is the value of vector ab is equal to???

Grade:12th pass

1 Answers

Arun
4 years ago

|AxB| = |A|*|B|*sin #

Where # is the angle between the two vectors.

Hence, sin # = |AxB| / (|A|*|B|) = 8/10

Hence, cos # = √(1-(8/10)²) = 6/10 or 3/5

|A|.|B| = |A|*|B|*cos # = 2*5*(3/5) = 6